OpenStudy (anonymous):

a²=x(10)^6 ÷ y/3 b²=x÷y given b = 8x10^5 find a

OpenStudy (anonymous):

anybody?

OpenStudy (phi):

dividing by y/3 is the same as multiplying by 3/y so you can write this as \[ a^2 = x \cdot 10^6 \cdot \frac{3}{y} \] we can re order the terms to look like this \[ a^2 = \frac{x}{y}\cdot 3 \cdot 10^6 \] notice you have x/y which you can replace with b^2 (the second equation tells you b^2 is the same as x/y) can you finish ?